OpenStudy (anonymous):

Solve x^2 = 14x – 49. Show Work

OpenStudy (anonymous):

\[x^2 -14x + 49 = 0\] you can factor \[(x-7)(x-7) = 0 \] so x = 7

hero (hero):

x^2-14x+49 = 0 x^2-7x-7x + 49 = 0 x(x-7)-7(x-7)=0 (x-7)(x-7)=0 (x-7)^2=0 x-7 =0 x = 7